OpenStudy (anonymous):

find the middle term of the expression(b/x+x/b)^16 using binomial theorem

OpenStudy (anonymous):

shouldnt it be 1? ([b/x]^8*[x/b]^8=1)

OpenStudy (anonymous):

the middle term = the 9th term =16 C 8 (b/x)^8 (x/b)^8= 12870 (1) = 12870